Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 995 hours each month to produce 1,990 sets of covers. The standard costs associated with this level of production are:

Total Per Set
of Covers
Direct materials $ 47,362 $ 23.80
Direct labor $ 8,955 4.50
Variable manufacturing overhead (based on direct labor-hours) $ 2,388 1.20
$ 29.50

During August, the factory worked only 1,000 direct labor-hours and produced 2,300 sets of covers. The following actual costs were recorded during the month:

Total Per Set
of Covers
Direct materials (8,800 yards) $ 50,600 $ 22.00
Direct labor $ 10,580 4.60
Variable manufacturing overhead $ 4,600 2.00
$ 28.60

At standard, each set of covers should require 3.5 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1
1. Standard price of materials per yard = Standard direct materials per set of covers / Standard quantity of materials per set of covers = 23.80 /3.5 6.8
Actual materials rate per yard = Actual total materials cost / Actual materials quantity = 50600 / 8800 5.75
Materials price variance = ( Actual price - Standard price ) * Actual quantity = ( 5.75 - 6.8 ) * 8800 9240 F
Standard quantity of materials for 2300 sets of covers = 2300 * Standard quantity of materials per set of covers = 2300 * 3.5 8050
Materials quantity variance = ( Actual quantity - Standard quantity ) * Standard price = ( 8800 - 8050 )* 6.8 5100 U
2.
Standard labor hours per sets of cover = Standard hours for 1990 sets of covers / 1990 = 995 / 1990 0.5
Standard rate of labor = Standard direct labor per set of cover / Standard labor hours per sets of covers = 4.50 / 0.5 9
Actual rate of labor = Actual total direct labor cost / Actual direct labor hours worked = 10580 / 1000 10.58
Labor rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 10.58 - 9 ) * 1000 1580 U
Standard direct labor hours for 2300 sets of covers = 2300 * Standard labor hours per sets of covers = 2300 * 0.5 1150
labor efficiency varaince = ( Actual hours - Standard hours ) * Standard rate = ( 1000 - 1150 ) * 9 1350 F
3.
Standard rate of variable overhead = Standard cost of variable overhead per sets of cover / Standard labor hours per sets of cover = 1.20 / 0.5 2.4
Actual rate of variable overhead = Actual total variable overhead cost / Actual direct labor hours worked = 4600 / 1000 4.6
Variable overhead rate variance = ( Actual rate - Standard rate ) * Actual hours = ( 4.6 - 2.4 ) * 1000 2200 U
Variable overhead efficiency variance = ( Actual hours - Standard hours ) * Standard rate = ( 1000 - 1150 ) * 2.4 360 F